# Creating General User Models from Computer Use

### v2

@@ Abstract.
Human-computer interaction has long imagined technology that understands us—from our preferences and habits, to the timing and purpose of our everyday actions. Yet current user models remain fragmented, narrowly tailored to specific applications, and incapable of the flexible, cross-context reasoning required to fulfill these visions. This paper presents an architecture for a general user model (GUM) that learns about you by observing any interaction you have with your computer. The GUM takes as input any unstructured observation of a user (e.g., device screenshots) and constructs confidence-weighted natural language propositions that capture that user’s behavior, knowledge, beliefs, and preferences. GUMs can infer that a user is preparing for a wedding they’re attending from a message thread with a friend. Or recognize that a user is struggling with a collaborator’s feedback on a draft paper by observing multiple stalled edits and a switch to reading related work. GUMs introduce an architecture that infers new propositions about a user from multimodal observations, retrieves related propositions for context, and continuously revises existing propositions. To illustrate the breadth of applications that GUMs enable, we demonstrate how they augment chat-based assistants with contextual understanding, manage OS notifications to surface important information only when needed, and enable interactive agents that adapt to user preferences across applications. We also instantiate a new class of proactive assistants (Gumbos) that discover and execute useful suggestions on a user’s behalf based on the their GUM. In our evaluations, we find that GUMs make calibrated and accurate inferences about users, and that assistants built on GUMs proactively identify and perform actions of meaningful value that users wouldn’t think to request explicitly. Altogether, GUMs introduce new methods that leverage large multimodal models to understand unstructured user context, enabling both long-standing visions of HCI and entirely new interactive systems that anticipate user needs. [^1] A demo of Gumbo and our open source package for GUM are available at https://generalusermodels.github.io
@@
copyright: noneccs: Computing methodologies Natural language processingccs: Human-centered computing Interactive systems and tools
@@ 1. Introduction
In this paper, we describe an architecture for general user models: computational models that materialize information and inferences about a user across domains and time scales. The general user model, or GUM, allows a user to construct a private, computational representation of their own behavior, knowledge, beliefs, and preferences by feeding unstructured observations (e.g., screenshares of their computer use) through an inference architecture leveraging large multimodal models (e.g., vision and language models, or VLMs). Our architecture contributes three main elements. A Propose module translates unstructured observations into confidence-weighted propositions about the user’s preferences, context, and intent. A Retrieve module indexes and searches these propositions to return the most contextually relevant subset for a given query. Finally, using results from Retrieve, a Revise module reevaluates and refines propositions as new observations arrive. We audit all observations for privacy violations with a contextual integrity (Nissenbaum, 2004) Audit module that leverages the GUM itself to estimate and filter out information that the user would not expect to be recorded into the GUM. All data stays securely on the user’s device, enabling local inference on capable hardware.
@@
In our technical evaluations, we first focus on validating GUM accuracy. We train GUM on recent email interaction, feeding each email—metadata, attachments, links, and replies—sequentially into the GUM. $N=18$ participants judged propositions generated by GUMs as overall accurate and well-calibrated: unconfident when incorrect, and confident when correct. Highly confident propositions (confidence = 10) were rated 100% accurate, while all propositions on average—including ones with low confidence—were fairly accurate (76.15%). From ablation studies, we show that all GUM components are critical for accuracy. We then deploy Gumbo with $N=5$ participants for 5 days, with the system observing the participants’ screens. This longitudinal evaluation replicated our results with the underlying GUM. Additionally, participants identified a meaningful number of useful and well-executed suggestions completed by Gumbo. Two of the five participants found particularly high value in the system and asked to continue running it on their computer after the study concluded. Our evaluations also highlight limitations and boundary conditions of GUM and Gumbo, including privacy considerations and overly candid propositions.
@@ 2. Related Work
The first major challenge that GUMs must engage with is common ground—the mutual knowledge, beliefs, and assumptions shared between individuals—which enables efficient communication (Clark and Schaefer, 1989). Through dialogue, gestures, facial expressions, eye gaze, and other multimodal cues (Clark, 1996), people continually update their shared mental models, iteratively constructing common ground (Clark and Brennan, 1991). Common ground acts as a shared context, allowing people to communicate without much unnecessary elaboration.
However, AI systems have little or zero shared context with users and make assumptions about a user’s background, producing one-size-fits-all answers that are over-informative (Tsvilodub et al., 2023), overconfident (Mielke et al., 2022), and unable to handle ambiguity (Abercrombie et al., 2023; Min et al., 2020; Gao et al., 2021). Inability to ground extends beyond question-asking. For example, designing and prototyping with autonomous AI systems necessitates collaborative grounding (Vaithilingam et al., 2024). Otherwise, they might take unfavorable and irreversible action (Bansal et al., 2024; Shao et al., 2024; Cemri et al., 2025).
To bridge the human-AI grounding gap, two general solutions have emerged: asking clarification questions or learning from a user’s dialogue history. Clarification-based grounding strategies attempt to simulate the natural back-and-forth dialogue that humans use to establish shared understanding. This interaction can be implemented through prompting (Kuhn et al., 2022; Chen et al., 2023), finetuning (Andukuri et al., 2024; Zhang and Choi, 2023; Hong et al., 2023; Gan et al., 2024), or a combination of both (Shaikh et al., 2025). Learning from or keeping a memory of a user’s chat interaction history is another solution (Lucas et al., 2009; Joshi et al., 2017; Zhang et al., 2018; Packer et al., 2023; Shaikh et al., [n. d.]). In practice, current models are limited to context derived solely from their own prior interactions with the user. All other computer interaction is inaccessible to models.
@@
The second major issue that GUMs draw on is that prompting LLMs effectively is difficult (Zamfirescu-Pereira et al., 2023b). Users often provide models with underspecified prompts (Zamfirescu-Pereira et al., 2023a; Shaikh et al., 2025; Agrawala, 2023), requiring repeated iteration to specify constraints for a specific context. A range of interactive systems have enabled users to better specify constraints to LLMs given a specific context: systems like ChainForge (Arawjo et al., 2024) and EvalGen (Shankar et al., 2024) offer users a means to interactively validate and iterate on prompts for specific tasks. Beyond developing prompts, systems like Graphologue (Jiang et al., 2023) and Sensescape (Suh et al., 2023) allow users to explore the range of possible outputs given a prompt. Across these interactions, the burden of providing explicit input to build a shared context is largely on the user.
@@
Still, our interactions with LLMs are far from visions like Weiser’s Sal, where computers unify machine intelligence with context from our daily lives (Horvitz, 2006). Interactive LLM-based systems focus on constraining the types of inputs, outputs, and tasks these models operate over. In this paper, we embrace what LLMs are good at—unconstrained input—and ground assumptions about the user and their context based on observation. We design what is effectively a multimodal grounding engine.
@@ 3. General User Models
In this section, we introduce an architecture for learning general user models. GUMs enable learning both stable inferences about the end user and moment-to-moment context, using everyday action as input data. From low-level input (e.g. screenshots), GUMs construct a rich understanding of who we are and what we are doing. Applications can then query the GUM to power a wide range of interactions, from operating systems that power contextual notifications to assistants that proactively discover and execute useful suggestions based on context.
@@
Proposition: [Anon] is currently writing in the General User Model’s section.
@@
Proposition: [Anon] is viewing and resolving comments from their advisor, [Anon].
@@
Proposition: [Anon] is struggling with the technical evaluation of GUMs.
@@
These propositions were generated entirely using a series of screenshots from the author’s screen. To construct propositions, GUM takes any type of input that an underlying large model can accept (in the case of vision-langauge models, GUM accepts image and text input). Using this unstructured input, GUM composes and refines proposition-confidence pairs. This abstraction—a series of confidence-weighed statements that conjecture some common ground between the user and their computer—enables a flexible representation of a user’s context. To be clear: GUMs do not enforce any structure on the types or kinds of propositions that can be generated. Propositions are constrained only by natural language descriptions, and are revised as the GUM continues receiving more input. GUMs aim to continually learn this representation of a user’s context from any interaction.
@@ 3.1. Propositions and Confidences: Abstractions for Representing General User Models
GUM draws inspiration from literature on common ground in human-human interaction (Clark and Brennan, 1991; Clark, 1996). A subset of common ground between participants can be defined by shared assumptions. The human-AI grounding challenge lies in identifying valid assumptions an AI model can make about a user. When constructing a user model, we must include shared assumptions that accurately represent user actions. For example, when a user repeatedly searches for ”best hiking trails in Colorado” and visits outdoor equipment websites, the model might form the proposition ”User enjoys outdoor activities, and is interested in hiking in Colorado” with high confidence. Similarly, if a user has numerous calendar entries for guitar lessons and frequently browses music theory websites, the model might infer ”User is learning to play guitar and is interested in music theory” as a shared assumption.
We build abstractions for GUM based on models of grounding in human-human communication (Brennan, 2014). Bayesian approaches represent shared knowledge as probabilistic distributions, not deterministic sets of propositions. In these models, interlocutors maintain probability distributions over possible states, refining a shared understanding through communication (Brennan, 2014). This probabilistic view accommodates user uncertainty and grounding over time (Lassiter, 2012; Horvitz et al., 2013). To instantiate a probabilistic model of common ground, we leverage the in-context learning capabilities of large langauge models, which can be loosely interpreted as Bayesian updates (Xie et al., 2021; Zhang et al., 2024a). We operationalize these approaches through three abstractions: observations, propositions, and confidences.
Observations are the raw inputs that the GUM processes. Observations can be anything, so long as it can be tokenized and fed into a large (vision/language/multimodal) model. Screenshot observations provide visual data of what the user is viewing, file system observations reveal document interactions, and notification observations offer communication pattern insights. Observations are the source from which GUM builds user context understanding. Each observation includes metadata: the source, timestamp, and a unique identifier. Unlike propositions, observations are factual records, not inferences. The relationship between observations and propositions is many-to-many. Multiple observations might support a single proposition, and a single observation might contribute to multiple propositions.
@@
Additionally, propositions include metadata for downstream applications. Inspired by cognitive agent architectures (Anderson, 1993; Park et al., 2023), we introduce generated decay factors (0-1), indicating proposition staleness rates. Author is a Ph.D. student decays slowly (factor = 1.0). Conversely, Author is debugging a memory leak is transient (factor = 0.6). Each proposition includes grounding: observations supporting the inference, plus generated reasoning explaining the proposition’s construction:
@@ 4.1. Prompting Models
Before we get distracted by visions, let’s return to something slightly more humble: prompting. A fundamental problem with prompting is a lack of shared context—called grounding (Shaikh et al., 2025)—with a model. Unless explicitly instructed, LLMs have no idea what project you’re working on, who you’re working with, or why the prompted task was important to you. This grounding gap results in models that often incorrectly assume aspects of our context.
@@
Proposition: [Anon] is currently focusing on the Applications section.
@@
Proposition: [Anon] recently read Weiser’s Calm Computing.
@@
Proposition: [Anon]’s advisor mentioning looking into Interface Agents.
@@ 4.2. Any Application Can Query A GUM
Applications of GUMs realize this vision. Operating systems that use GUMs can surface important information only when it’s truly needed, avoiding information overload. For example, the first author of this paper was swamped with writing this paper, and the CHI early bird registration deadline came and went. By default, their operating system shows notifications for every email they gets in their primary inbox—causing them to generally ignore the notification, as they also received emails in the same period about “Ice Cream Recipe Found!”, “firebase-noreply-billing”, and “Job Opportunity at Startup”. When prompting Llama 3.3 with “Filter out notifications that are important to me.”, it suggests passing all of these notifications through to the author’s attention, leading to our poor author’s attention overload. However, the same prompt, augmented with GUM’s propositions and confidence, provides context that the author (1) “is quite happy as a Ph.D. student”, (2) “has a major impending deadline”, (3) “regularly ignores billing notifications”, and (4) “makes ice-cream on weekends as a hobby.” As a result, the LLM now allows only the CHI registration notification through.
@@
Interface agents introduced personalized and proactive assistance as a vision for human-computer interaction. Instead of placing the burden of initiative on the user, interface agents are applications that aim to anticipate user needs and automate tasks. This requires both domain knowledge and personalized user models, but an implementation has remained challenging without a continuously learned model of individual preferences. Each new agent requires specialized datasets and hand-crafted learning objectives to capture evolving user preferences (Maes and Kozierok, 1993).
@@
Proposition: [Anon] is prioritizing their paper deadline over routine physical exercise.

### v3

@@ Abstract.
Human-computer interaction has long imagined technology that understands us—from our preferences and habits, to the timing and purpose of our everyday actions. Yet current user models remain fragmented, narrowly tailored to specific applications, and incapable of the flexible, cross-context reasoning required to fulfill these visions. This paper presents an architecture for a general user model (GUM) that learns about you by observing any interaction you have with your computer. The GUM takes as input any unstructured observation of a user (e.g., device screenshots) and constructs confidence-weighted natural language propositions that capture that user’s behavior, knowledge, beliefs, and preferences. GUMs can infer that a user is preparing for a wedding they’re attending from a message thread with a friend. Or recognize that a user is struggling with a collaborator’s feedback on a draft paper by observing multiple stalled edits and a switch to reading related work. GUMs introduce an architecture that infers new propositions about a user from multimodal observations, retrieves related propositions for context, and continuously revises existing propositions. To illustrate the breadth of applications that GUMs enable, we demonstrate how they augment chat-based assistants with contextual understanding, manage OS notifications to surface important information only when needed, and enable interactive agents that adapt to user preferences across applications. We also instantiate a new class of proactive assistants (Gumbos) that discover and execute useful suggestions on a user’s behalf based on their GUM. In our evaluations, we find that GUMs make calibrated and accurate inferences about users, and that assistants built on GUMs proactively identify and perform actions of meaningful value that users wouldn’t think to request explicitly. Altogether, GUMs introduce new methods that leverage large multimodal models to understand unstructured user context, enabling both long-standing visions of HCI and entirely new interactive systems that anticipate user needs.
@@
journalyear: 2025copyright: acmlicensedconference: The 38th Annual ACM Symposium on User Interface Software and Technology; September 28-October 1, 2025; Busan, Republic of Koreabooktitle: The 38th Annual ACM Symposium on User Interface Software and Technology (UIST ’25), September 28-October 1, 2025, Busan, Republic of Koreadoi: 10.1145/3746059.3747722isbn: 979-8-4007-2037-6/2025/09ccs: Computing methodologies Natural language processingccs: Human-centered computing Interactive systems and tools
@@ 1. Introduction
In this paper, we describe an architecture for general user models: computational models that materialize information and inferences about a user across domains and time scales. [^1] A demo of Gumbo and our open source package for GUM are available at https://generalusermodels.github.io The general user model, or GUM, allows a user to construct a private, computational representation of their own behavior, knowledge, beliefs, and preferences by feeding unstructured observations (e.g., screenshares of their computer use) through an inference architecture leveraging large multimodal models (e.g., vision and language models, or VLMs). Our architecture contributes three main elements. A Propose module translates unstructured observations into confidence-weighted propositions about the user’s preferences, context, and intent. A Retrieve module indexes and searches these propositions to return the most contextually relevant subset for a given query. Finally, using results from Retrieve, a Revise module reevaluates and refines propositions as new observations arrive. We audit all observations for privacy violations with a contextual integrity (Nissenbaum, 2004) Audit module that leverages the GUM itself to estimate and filter out information that the user would not expect to be recorded into the GUM. All data stays securely on the user’s device, enabling local inference on capable hardware.
@@
In our technical evaluations, we first focus on validating GUM accuracy. We train GUM on recent email interaction, feeding each email—metadata, attachments, links, and replies—sequentially into the GUM. Emails contain diverse and relevant information across modalities, while also being full of irrelevant content. This makes synthesizing important information difficult: GUMs must maintain accuracy despite this challenge. Still, $N=18$ participants judged propositions generated by GUMs as overall accurate and well-calibrated: unconfident when incorrect, and confident when correct. Highly confident propositions (confidence = 10) were rated 100% accurate, while all propositions on average—including ones with low confidence—were fairly accurate (76.15%). From ablation studies, we show that all GUM components are critical for accuracy. We then deploy Gumbo with $N=5$ participants for 5 days, with the system observing the participants’ screens. This longitudinal evaluation replicated our results with the underlying GUM. Additionally, participants identified a meaningful number of useful and well-executed suggestions completed by Gumbo. Two of the five participants found particularly high value in the system and asked to continue running it on their computer after the study concluded. Our evaluations also highlight limitations and boundary conditions of GUM and Gumbo, including privacy considerations and overly candid propositions.
@@ 2. Related Work
The first major challenge that GUMs must engage with is common ground: the mutual knowledge, beliefs, and assumptions shared between individuals that enables efficient communication (Clark and Schaefer, 1989). Through dialogue, gestures, facial expressions, and other multimodal cues (Clark, 1996), people continually update shared mental models, iteratively constructing common ground (Clark and Brennan, 1991). Common ground acts as a shared context, allowing people to communicate without unnecessary elaboration.
However, AI systems have little shared context with users and make assumptions about a user’s background, producing one-size-fits-all answers that are over-informative (Tsvilodub et al., 2023), overconfident (Mielke et al., 2022), and unable to handle ambiguity (Abercrombie et al., 2023; Min et al., 2020; Gao et al., 2021). Inability to ground extends beyond question-asking. For example, designing and prototyping AI systems requires collaborative grounding (Vaithilingam et al., 2024). Otherwise, they might take unfavorable and irreversible action (Bansal et al., 2024; Shao et al., 2024; Cemri et al., 2025).
To bridge the human-AI grounding gap, two general solutions have emerged: asking clarification questions or learning from a user’s dialogue history. Clarification-based grounding attempts to simulate the natural back-and-forth dialogue that humans use to establish shared understanding. This interaction can be implemented through prompting (Kuhn et al., 2022; Chen et al., 2023), finetuning (Andukuri et al., 2024; Zhang and Choi, 2023; Hong et al., 2023; Gan et al., 2024), or a combination of both (Shaikh et al., 2025). Keeping a long-term memory of a user’s chat interaction history is another solution (Lucas et al., 2009; Joshi et al., 2017; Zhang et al., 2018; Packer et al., 2023; Shaikh et al., [n. d.]). Long-context models also allow users to provide lengthy inputs as grounding (Team et al., 2024), but models generally forget information as context lengths increase (Liu et al., 2023). Meanwhile, retrieval-augmented generation (RAG) (Lewis et al., 2020)—where context is retrieved from a traditional database—has served as an alternative approach. For example, systems like OmniQuery (Li et al., 2025) draw from a processed database of multimodal memories from photo albums, contextualizing queries to an LLM. Regardless of the approach, current models are limited to context derived solely from their own prior interactions with the user or from a limited taxonomy of pre-defined applications or sources. All other computer interaction is inaccessible to models.
@@
The second major issue that GUMs draw on is that prompting LLMs effectively is difficult (Zamfirescu-Pereira et al., 2023b). Users often provide models with underspecified prompts (Zamfirescu-Pereira et al., 2023a; Shaikh et al., 2025; Agrawala, 2023), requiring repeated iteration to specify constraints for a specific context. A range of interactive systems have enabled users to better specify constraints to LLMs: systems like ChainForge (Arawjo et al., 2024) and EvalGen (Shankar et al., 2024) offer users a means to interactively validate and iterate on prompts for specific tasks. Beyond developing prompts, systems like Graphologue (Jiang et al., 2023) and Sensescape (Suh et al., 2023) allow users to explore the range of possible outputs given a prompt. Across these interactions, the burden of providing explicit input to build shared context lies with the user.
@@
Still, our interactions with LLMs are far from visions like Weiser’s Sal, where computers unify machine intelligence with context from our daily lives (Horvitz, 2006). Instead, LLM-based systems focus on constraining the types of inputs, outputs, and tasks models operate over. In this paper, we embrace what LLMs are good at—unconstrained input—and ground assumptions about the user and their context based on observation. We design what is effectively a multimodal grounding engine.
@@ 3. General User Models
We introduce an architecture for learning general user models. GUMs enable learning both stable inferences about the end user and moment-to-moment context, using everyday action as input data. From low-level input (e.g. screenshots), GUMs construct a rich understanding of who we are and what we are doing. Applications can then query the GUM to power a range of interactions, from operating systems that power contextual notifications to assistants that proactively discover and execute useful suggestions based on context.
@@
Proposition: Omar is currently writing in the General User Model’s section.
@@
Proposition: Omar is viewing and resolving comments from their advisors, Diyi and Michael.
@@
Proposition: Omar is struggling with the technical evaluation of GUMs.
@@
Propositions were generated entirely using a series of screenshots from the author’s screen. GUM takes any type of input that an underlying large model can accept (in the case of vision-langauge models, GUM accepts image and text input). Using this unstructured input, GUM composes and refines proposition-confidence pairs. This abstraction—a series of confidence-weighed statements that conjecture some common ground between the user and their computer—enables a flexible representation of a user’s context. To be clear: GUMs do not enforce any structure on the types or kinds of propositions that can be generated. Propositions are constrained only by natural language descriptions, and are revised as the GUM continues receiving more input. GUMs aim to continually learn this representation of a user’s context from any interaction.
@@ 3.1. Propositions and Confidences: Abstractions for Representing General User Models
GUM draws inspiration from literature on common ground in human-human interaction (Clark and Brennan, 1991; Clark, 1996). A subset of common ground between participants can be defined by shared assumptions. The human-AI grounding challenge lies in identifying valid assumptions an AI model can make about a user. When constructing a user model, we must include shared assumptions that accurately represent user actions. For example, when a user repeatedly searches for ”best hiking trails in Colorado” and visits outdoor equipment websites, the model might form the proposition ”User enjoys outdoor activities, and is interested in hiking in Colorado” with high confidence.
We build abstractions for GUM based on grounding models in human-human communication (Brennan, 2014). Bayesian approaches represent shared knowledge as probabilistic distributions, not deterministic sets of propositions. In these models, interlocutors maintain probability distributions over possible states, refining a shared understanding through communication (Paek and Horvitz, 1999, 2013; Goodman and Frank, 2016). This probabilistic view accommodates user uncertainty and grounding over time (Lassiter, 2012; Horvitz et al., 2013). To instantiate a probabilistic model of common ground, we leverage the in-context learning capabilities of large langauge models, which can be loosely interpreted as Bayesian updates (Xie et al., 2021; Zhang et al., 2024a). We operationalize these approaches through three abstractions: observations, propositions, and confidences.
Observations are raw inputs that the GUM processes. Observations can be anything, so long as they can be tokenized and fed into a large (vision/language/multimodal) model. Screenshot observations provide visual data of what the user is viewing, file system observations reveal document interactions, and notification observations offer communication pattern insights. Observations are the source from which GUM builds user context understanding. Each observation includes metadata: the source, timestamp, and a unique identifier. Unlike propositions, observations are factual records, not inferences.
@@
Additionally, propositions include metadata for downstream applications. Inspired by cognitive agent architectures (Anderson, 1993; Park et al., 2023), we introduce generated decay scores (0-1), indicating proposition staleness rates. Author is a Ph.D. student decays slowly (score = 1.0). Conversely, Author is debugging a memory leak is transient (score = 0.6). Each proposition includes grounding: observations supporting the inference, plus generated reasoning explaining the proposition’s construction:
@@ 4.1. Prompting Models
Before we get distracted by visions, let’s return to something slightly more humble: prompting. A fundamental problem with prompting is a lack of shared grounding (Shaikh et al., 2025) with a model. Unless explicitly instructed, LLMs have no idea what project you’re working on, who you’re working with, or why the prompted task was important to you. This grounding gap results in models that often incorrectly assume aspects of our context.
@@
Proposition: Omar is currently focusing on the Applications section.
@@
Proposition: Omar recently read Weiser’s Calm Computing.
@@
Proposition: Omar’s advisor mentions looking into Interface Agents.
@@ 4.2. Any Application Can Query A GUM
For example, operating systems that use GUMs can surface important information when it’s truly needed. For example, the first author of this paper was swamped with writing this paper, and the CHI early bird registration deadline came and went. By default, their operating system shows notifications for every email they gets in their primary inbox—causing them to generally ignore the notification, as they also received emails in the same period about “firebase-noreply-billing” and “Job Opportunity at Startup”. When prompting Llama 3.3 with “Filter out notifications that are important to me.”, it suggests passing all of these notifications through to the author, leading to attention overload. However, the same prompt, augmented with GUM’s propositions and confidence, provides context that the author (1) “is quite happy as a Ph.D. student”, (2) “has a major impending deadline”, and (3) “regularly ignores billing notifications” As a result, the LLM now allows only the CHI registration notification through.
@@
Interface agents introduced personalized and proactive assistance for human-computer interaction. Instead of placing the burden of initiative on the user, interface agents are applications that aim to anticipate user needs and automate tasks. This requires both domain knowledge and personalized user models, but an implementation has remained challenging without a continuously learned model of individual preferences. Each new agent requires specialized datasets and hand-crafted learning objectives to capture evolving user preferences (Maes and Kozierok, 1993).
@@
Proposition: Omar is prioritizing their paper deadline over routine physical exercise.